OpenStudy (charlotte123):

Please help! :) Gino designed a garden for his rectangular backyard. The design he drew on paper is 7 inches wide. The width of his backyard is 42 feet. If his drawing is to scale and is 8 inches long, how many feet long is his backyard? Write a proportion using words. Then, use the proportion to calculate the yard length. Show your work.

OpenStudy (zzr0ck3r):

42/7=x/8 x = 8(42/7) I think

OpenStudy (anonymous):

@zzr0ck3r what he said

OpenStudy (charlotte123):

thanks for the help everyone :D

OpenStudy (charlotte123):

but what does x represent?

OpenStudy (zzr0ck3r):

length of hte yard

OpenStudy (charlotte123):

thank u :)

OpenStudy (zzr0ck3r):

you could have said 7/42 = 8/x x(7/42) = 8 x = (42/7)8

OpenStudy (zzr0ck3r):

this is the cool thing about ratios:)

OpenStudy (charlotte123):

:D so x = 48? :)

OpenStudy (zzr0ck3r):

nono x = (42/7)*8

OpenStudy (charlotte123):

which is 48? :P

OpenStudy (zzr0ck3r):

o lol yeah sorry i thought you wrote 42

OpenStudy (charlotte123):

its ok :P so 42/7 = x/8 x = (42/7) * 8 x = 48? :)

OpenStudy (zzr0ck3r):

looks good to me

OpenStudy (charlotte123):

thanks for all the help :D have a great day ;)

OpenStudy (charlotte123):

wait but it said to write a proportion using words :O

OpenStudy (charlotte123):

so the proportion in words is? sorry just confused :(

OpenStudy (zzr0ck3r):

width of yard over width of model = length of yard over length of model

OpenStudy (charlotte123):

u mean Width of yard over width of design = length of yard over length of drawing?

OpenStudy (zzr0ck3r):

yeah

OpenStudy (charlotte123):

ok thanks :)